OpenStudy (anonymous):

Fimd the volume of the cylinder. Round to the nearest tenth of a square unit. d=28.6in., h=10.7in

OpenStudy (anonymous):

v = πr^2 * h

OpenStudy (anonymous):

substitute in for r = 1/2d, h = 10.7

OpenStudy (anonymous):

actually its cubic units for volume

OpenStudy (anonymous):

v=3.14x14.3x14.3x10.7=6870.45502

OpenStudy (anonymous):

6873.93982 in^3 and then rounded 6874.0 in^3 Rounding π early may skew the answer

OpenStudy (anonymous):

what is skew? I'm sorry I haven't been in a math class in 15 yrs. Is that the answer ur last post

OpenStudy (anonymous):

Skew as in change the result, if you don't round π to 3.14 and you use 3.141592653589793 you get the result I posted: 6874.0 in^3

OpenStudy (anonymous):

ok thank u very much, this math is kicking my behind. LOL

OpenStudy (anonymous):

V =pi r^2 h = 3.14 *14.3^2 *10.7 = 6.9